What mass of KF must be added to 200 mL of 0.15 M HF to obtain a buffer with pH = 3.00? Xg Enter a number.a)What is the pH of a buffer made by mixing 75 mL of 0.1 M hydrocyanic acid (HCN) with 100 mL of 0.1 M NaCN at 25 °C? 9.43 b)How many grams of solid NaCNO must be added to 100 mL of 0.2 M HCNO at 25°C to obtain a buffer at pH = 4.00?
Answer #1

1)

pH = 3.00

pKa = 3.20

moles of HF = 200 x 0.15 / 1000 = 0.030 mol

pH = pKa + log [salt / acid]

3.00 = 3.20 + log [KF / 0.030]

moles of KF = 0.0189

mass of KF added = 0.0189 x 58

mass of KF added = 1.10 g

2)

mmoles of HCN = 75 x 0.1 = 7.5

mmoles of NaCN = 100 x 0.1 = 10

mmoles of NaOH = 2 x 1 = 2

pH = pKa + log [salt + C / acid - C]

     =9.305 + log [10 + 2 / 7.5 - 2]

pH = 9.64

3)

pH = pKa + log [NaCNO / HCNO]

4.00 = 3.456 + log [NaCNO / 0.1 x 0.2]

moles of NaCNO = 0.070

mass of NaCNO = 0.07 x 65

                            = 4.55 g

Note : I took the pKa values from the appendix which i have. please check once whether these match with your values or not. if any doubt ask me. if any wrong that is only with pKa values